a mercury barometer reads the atmospheric pressure on top of a mountain as 74.5cmHg and the base of the mountain it reads 76cmHg. what is the height of the mountain


1
Expert's answer
2021-03-14T19:17:54-0400

It is known, that for small altitudes the atmospheric pressure decreases by 1 mmHG for each 12 m of raising. So, the height of the mountain is:


"h=12(p_b-p_t),"



where pt - the atmospheric pressure on top of a mountain in mmHg;

pb - the atmospheric pressure on base of a mountain in mmHg.


"h=12(760-745)=180\space m."


Answer: 180 m.